# LV-XAttn: Distributed Cross-Attention for Long Visual Inputs in Multimodal Large Language Models

### v1

@@ Abstract
Cross-attention is commonly adopted in multimodal large language models (MLLMs) for integrating visual information into the language backbone. However, in applications with large visual inputs, such as video understanding, processing a large number of visual tokens in cross-attention layers leads to high memory demands and often necessitates distributed computation across multiple GPUs. Existing distributed attention mechanisms face significant communication overheads, making cross-attention layers a critical bottleneck for efficient training and inference of MLLMs. To address this, we propose LV-XAttn, a distributed, exact cross-attention mechanism with minimal communication overhead. We observe that in applications involving large visual inputs the size of the query block is typically much smaller than that of the key-value blocks. Thus, in LV-XAttn we keep the large key-value blocks locally on each GPU and exchange smaller query blocks across GPUs. We also introduce an efficient activation recomputation technique enabling support for longer visual context. We theoretically analyze the communication benefits of LV-XAttn and show that it can achieve speedups for a wide range of models. Our evaluations with mPLUG-Owl3 and OpenFlamingo models find that LV-XAttn achieves up to 5.58$\times$ end-to-end speedup compared to existing approaches.
@@ 1 Introduction
A common approach to integrating visual inputs into LLMs is through cross-attention (Alayrac et al., 2022; Laurençon et al., 2024; Li et al., 2023a; Ye et al., 2024; Grattafiori et al., 2024), where queries derived from the text input interact with keys and values derived from the visual inputs. This enables effective fusion of multimodal information. In MLLMs, cross-attention layers are inserted between language model blocks, enabling the LLM to process intermediate representations that are integrated with visual information.
[…]
@@
Existing distributed attention approaches can be categorized as two classes: head-parallelism and sequence-parallelism. Head-parallelism methods such as Deepspeed-Ulysses (Jacobs et al., 2024) and Megatron-LM (Korthikanti et al., 2023a) partition the computation along the head dimension of multi-head attention. Consequently, maximum degree of parallelism is capped by the number of heads used in multi-head attention. This translates to an upper bound in terms of memory capacity, preventing them from processing longer visual inputs which have memory demands beyond this (Table 6 in Section 4). In addition, the number of workers must be divisible by the total number of heads to ensure a balanced load across workers. Otherwise, resource underutilization may occur due to stragglers. On the other hand, sequence parallel methods such as Ring Attention (Liu et al., 2024a) partition the computation along the input sequence dimension, overcoming the limitation of head-parallelism methods. However, when applied to cross-attention with large visual inputs, these approaches suffer from large communication overheads even after overlapping computation and communication. Figure 2 shows that cross-attention operations distributed with Ring Attention (Liu et al., 2024a) can account for up to 87% of the iteration time, despite comprising only 2% of the total parameters.
@@
In this work, we present LV-XAttn, a distributed, exact cross-attention mechanism that employs sequence-parallelism with minimal communication overhead. Our main observation is that while keys and values derived from visual inputs are large, the queries derived from text input are typically small in MLLMs. For example, in the video understanding benchmark Video-MME (Fu et al., 2024), an input processed with mPLUG-Owl3 models results in an average sequence length of 1,739,394 for keys and values and 5,514 for queries, when frames are sampled at 1 fps. Based on this, LV-XAttn organizes each worker to locally store a partition of the large key and value blocks, while small query blocks are transmitted between workers to compute the attention output in a blockwise fashion. This significantly reduces the communication volume compared to Ring Attention. For instance, with the Video-MME benchmark, LV-XAttn reduces communication volume to just 0.48% of that required by Ring Attention. Furthermore, the reduced communication can be effectively overlapped by computation, allowing distributed cross-attention to be performed without incurring any communication overhead.
@@
To further enable the processing of longer visual inputs, we employ an activation recomputation technique that is specific to MLLMs. In standard attention implementations, activations including queries, keys, and values need to be saved for backward pass (Korthikanti et al., 2023b). Storing the large key and value tensors for every cross-attention layer introduces additional memory pressure. We observe that since cross-attentions in MLLMs share input visual tokens, we can maintain a single copy of the visual tokens accessible to all cross-attention layers and recompute activations during the backward pass. This allows us to process up to 1.6$\times$ longer visual inputs with just less than 8% overhead.
@@
We perform comprehensive evaluation of LV-XAttn on mPLUG-Owl3 models and OpenFlamingo (Awadalla et al., 2023) models across multiple cluster configurations, including setups with A100 and A30 GPUs. LV-XAttn speeds up the cross-attention operation by up to 45.85$\times$ and overall model iteration time by up to 5.58$\times$ compared to Ring Attention. By minimizing communication volume and further overlapping communication with computation, we demonstrate that LV-XAttn incurs less than 0.42% overhead compared to a no-communication baseline.
@@ 2 Background
Multimodal Large Language Models As LLMs continue to evolve, researchers are investigating how to incorporate vision and other modalities into these models. One common way is to embed cross-attention layers into the language model. This design has been adopted by a number of models including Flamingo (Alayrac et al., 2022), Otter (Li et al., 2023a), mPLUG-Owl3 (Ye et al., 2024), IDEFICS (Laurençon et al., 2024), and LLama 3-V (Grattafiori et al., 2024). Broadly, these models follow the architecture illustrated in Figure 1. They include a visual encoder, a projection layer, and an LLM. Cross-attention layers are interleaved between the layers of the LLM. Language inputs are fed directly into the LLM and the resulting intermediate representations are passed to the cross-attention layers as $x$. Visual inputs are processed by the visual encoder and the projection layer to produce visual tokens, which are then passed to the cross-attention layers as $y$, enabling the incorporation of visual information.
@@
Figure 2: Runtime breakdown for a single iteration of mPLUG-Owl-7b and OpenFlamingo-3b with Ring Attention and LV-XAttn on 16 A100 GPUs. mPLUG-Owl-7b has 4 cross-attention layers distributed across its 28 LM blocks, while OpenFlamingo-3b inserts 1 cross-attention layer after each of its 24 LM blocks. LV-XAttn reduces the time spent on cross-attention computation by 93% for mPLUG-Owl-7b and by 53% for OpenFlamingo-3b, compared to Ring Attention.
When applied to scenarios with large visual inputs, cross-attention requires significant memory resources and hence present a scaling challenge. The standard implementation of attention involves materializing the matrix product $QK^{T}\in\mathbb{R}^{S_{Q}\times S_{KV}}$, resulting in memory complexity that scales with the product of text sequence and the visual sequence length. The large amount of visual tokens from long videos inputs thus causes a memory bottleneck. For example, in mPLUG-Owl3, an hour-long video sampled at 1 fps is encoded to a visual input with more than 2 million tokens.
While memory-efficient methods like Flash Attention (Dao et al., 2022; Dao, 2024) reduce the memory footprint of attention operations to enable handling longer context lengths, the amount of memory required still often surpasses the capacity of a single worker. For example, processing an hour-long video with a language input of 4096 tokens in mPLUG-Owl3 demands 298 GB of memory for the cross-attention operation, even with Flash Attention.
@@
$O_{i}$ is computed iteratively by transmitting key-value blocks among workers in a ring-like fashion. To facilitate the block-wise computation of $O_{i}$, worker $i$ has to maintain necessary softmax statistics $m_{i}\in\mathbb{R}^{\frac{S_{Q}}{n}\times d}$ and $l_{i}\in\mathbb{R}^{\frac{S_{Q}}{n}}$. During round $r$, worker $i$ computes partial attention using the blocks $Q_{i}$, $K_{(i-r)\mod n}$, and $V_{(i-r)\mod n}$ and updates $O_{i},m_{i}$ and $l_{i}$. The key-value block is then sent to worker $i+1$ while a new key-value block is received from worker $i-1$.
While Ring Attention can be used to distribute the cross-attention operation, the presence of large key-value blocks makes Ring Attention communication-bound, resulting in highly inefficient cross-attention operations. As illustrated in Figure 2, despite comprising only 2% of the total parameters, cross-attention operations can account for up to 87% of the iteration time when using Ring Attention.
@@ 3.1 Method
         Send $O_{j_{\text{prev}}},m_{j_{\text{prev}}},l_{j_{\text{prev}}},Q_{j}$ to worker $(i+1)\bmod n$
         Recv $O_{j},m_{j},l_{j},Q_{j_{\text{next}}}$ from worker $(i-1)\bmod n$
         $\Delta m,\Delta l,\Delta O\leftarrow\text{FlashAttention}(Q_{j},K_{i},V_{i})$
     $O_{j},m_{j},l_{j}\leftarrow\text{Rescale}(O_{j},m_{j},l_{j},\Delta O,\Delta m,\Delta l)$
@@
The primary observation that motivates our work is that in applications involving large visual inputs, the size of the query block is typically much smaller than that of the key-value blocks. For instance, in the widely-used video understanding benchmark, Video-MME (Fu et al., 2024), the average text prompt for long videos consists of $S_{Q}=3128$ words, including the question, options, answer, and subtitles. On the other hand, videos have an average duration of 2,386 seconds; each frame is encoded by the visual encoder and projection layer in an MLLM into multiple visual tokens. For example, mPLUGOwl-3 generates 729 visual tokens per frame. With a sampling rate of 1 fps, each video results in a sequence length of $S_{KV}=1739394$. As a result, distributed attention mechanisms that involves movement of key-value blocks incur substantial communication overhead.
To address this, we propose LV-XAttn, which keeps the large key-value blocks locally on each worker, while smaller query blocks, attention blocks, and necessary softmax statistics are exchanged among workers in a ring-style fashion. This is illustrated in Figure 3. During each round, each worker $i$ computes attention using its local key-value blocks $K_{i}$ and $V_{i}$ and query blocks $Q_{j}$ received from peers. This computation generates partial attention blocks $\Delta O$ and partial softmax statistics $\Delta m$ and $\Delta l$. The worker then updates the received attention block $O_{j}$ and softmax statistics $m_{j}$ and $l_{j}$ by rescaling them using $\Delta O$, $\Delta m$, and $\Delta l$. The worker then sends $Q_{j}$, $O_{j}$, $m_{j}$, and $l_{j}$ to the next worker in the ring topology and receives $Q_{j-1}$, $O_{j-1}$, $m_{j-1}$, and $l_{j-1}$ from the previous worker. After $n$ rounds, the computed attention block $O_{i}$ and softmax statistics $m_{i}$ and $l_{i}$ are returned to worker $i$.
@@
  • Receive $O_{j}$, $m_{j}$ and $l_{j}$ from worker $i-1$, which are needed for rescaling in this round.
@@
  • Send $O_{j+1}$, $m_{j+1}$, $l_{j+1}$ computed in the previous round to worker $i+1$.
@@
After receiving $O_{j}$, $m_{j}$ and $l_{j}$, and computing $\Delta O,\Delta m$ and $\Delta l$, we can perform rescaling to update $O_{j}$, $m_{j}$ and $l_{j}$. We describe our distributed attention procedure in Algorithm 1. As demonstrated in Section 4.4, the substantial reduction in communication volume enables complete overlap with computation, effectively eliminating any communication overhead.
Figure 4: The theoretical speedup of LV-XAttn over Ring Attention for cross-attention on a 4-node cluster. Each node is equipped with 4 A100 GPUs, and nodes are interconnected by a 25 GB/s network. The markers represent processing a 2,386-second video and a 3,128-word text prompt—average values for long videos in Video-MME (Fu et al., 2024)—using mPLUG-Owl3 and OpenFlamingo models at different frame rates. Note that for each frame, a special token <image> have to be added to the text-prompt, resulting in $S_{Q}=2386+3128=5514$.
Table 1: Runtime analysis for Ring Attention and LV-XAttn. $h$ represents the number of head in multi-head attention. The attention FLOPs are calculated similar to previous work (Dao, 2024).
@@
$\max\left(f(\frac{S_{Q}d}{n},\frac{S_{KV}d}{n}),comm(3\cdot\frac{S_{Q}d}{n}+\frac{S_{Q}}{n})\right)$ (1)
@@ 3.2 Activation Recomputation for MLLM
In standard attention implementation, during forward pass computation, input tensors $Q_{i},K_{i},V_{i}$ and output tensors $O_{i}$ and $L_{i}$ are saved for backward pass, where $L_{i}=m_{i}+\log{l_{i}}$ (Dao, 2024). However, storing large key-value blocks $K_{i}$ and $V_{i}$ increases memory usage, thereby limiting the maximum number of visual inputs that can be processed. For instance, in the case of mPLUG-Owl3-7b with a 4096-frame video, storing $K_{i}$ and $V_{i}$ takes $79.73$GB per cross-attention layer.
@@
As demonstrated in the ablation study in Section 4.4, this approach incurs an overhead of less than $8\%$ while enabling the system to handle 1.6×1.6$\times$1.6 × more visual inputs.

### v3

@@ Abstract
Cross-attention is commonly adopted in multimodal large language models (MLLMs) for integrating visual information into the language backbone. However, in applications with large visual inputs, such as video understanding, processing a large number of visual tokens in cross-attention layers leads to high memory demands and often necessitates distributed computation across multiple GPUs. Existing distributed attention mechanisms face significant communication overheads, making cross-attention layers a critical bottleneck for efficient training and inference of MLLMs. To address this, we propose LV-XAttn, a distributed, exact cross-attention mechanism with minimal communication overhead. We observe that in applications involving large visual inputs, the size of the query block is typically much smaller than that of the key-value blocks. Thus, in LV-XAttn we keep the large key-value blocks locally on each GPU and exchange smaller query blocks across GPUs. We also introduce an efficient activation recomputation technique to support longer visual context. We theoretically analyze the communication benefits of LV-XAttn and show that it can achieve speedups for a wide range of models. Our evaluations with Llama 3-V, mPLUG-Owl3 and OpenFlamingo models find that LV-XAttn achieves up to 10.62$\times$ end-to-end speedup compared to existing approaches.
@@ 1 Introduction
A common approach to integrating visual inputs into LLMs is through cross-attention (Alayrac et al., 2022; Laurençon et al., 2024; Li et al., 2023a; Ye et al., 2024; Grattafiori et al., 2024; Dai et al., 2024), where queries derived from the text input interact with keys and values derived from the visual inputs. This enables effective fusion of multi-modal information. In MLLMs, cross-attention layers are inserted between language model blocks, enabling the LLM to process intermediate representations that are integrated with visual information.
However, the memory requirement of cross-attention layers is a limiting factor for applications involving large visual inputs, such as long video understanding. For example, in Llama 3-V (Grattafiori et al., 2024), applying cross-attention to a text sequence of length 2048 and a 20-minute video sampled at 1 frame per second (fps) requires over 234 GB of memory, even with a memory-efficient attention implementation (Dao, 2024). This exceeds the memory capacity of existing accelerators, necessitating the distributed computation of the attention operation across multiple workers.
@@
Existing distributed attention approaches can be categorized into two classes: head-parallelism and sequence-parallelism. Head-parallelism methods such as Deepspeed-Ulysses (Jacobs et al., 2024) and Megatron-LM (Korthikanti et al., 2023a) partition the computation along the head dimension of multi-head attention. Consequently, maximum degree of parallelism is capped by the number of heads used in multi-head attention. This translates to an upper bound in terms of memory capacity, preventing them from processing longer visual inputs which have memory demands beyond this. In addition, the number of workers must be divisible by the total number of heads to ensure a balanced load across workers. Otherwise, resource underutilization may occur due to stragglers. On the other hand, sequence parallel methods such as Ring Attention (Liu et al., 2024a) partition the computation along the input sequence dimension, overcoming the limitation of head-parallelism methods. However, when applied to cross-attention with large visual inputs, these approaches suffer from large communication overheads even after overlapping computation and communication. Figure 2 shows that cross-attention operations distributed with Ring Attention can account for up to 88% of the iteration time, despite comprising only 3% of the total parameters.
@@
In this work, we present LV-XAttn, a distributed, exact cross-attention mechanism that employs sequence-parallelism with minimal communication overhead. Our main observation is that while keys and values derived from visual inputs are large, the queries derived from text input are typically small in MLLMs. For example, in the video understanding benchmark Video-MME (Fu et al., 2024), an input processed with Llama 3-V results in an average sequence length of 15,279,944 for keys and values and 5,514 for queries, when frames are sampled at 1 fps. Based on this, LV-XAttn organizes each worker to locally store a partition of the large key and value blocks, while small query blocks are transmitted between workers to compute the attention output in a blockwise fashion. This significantly reduces the communication volume compared to Ring Attention. For instance, on the Video-MME benchmark, LV-XAttn reduces communication volume to just 0.04% of that required by Ring Attention. Furthermore, the reduced communication can be effectively overlapped by computation, allowing distributed cross-attention to be performed without incurring any communication overhead.
@@
To further enable the processing of longer visual inputs, we employ an activation recomputation technique that is specific to MLLMs. In standard attention implementations, activations including queries, keys, and values need to be saved for backward pass (Korthikanti et al., 2023b). Storing the large key and value tensors for every cross-attention layer introduces additional memory pressure. We observe that since cross-attention layers in MLLMs share the same input visual tokens, we can maintain a single copy of the visual tokens accessible to all cross-attention layers and recompute activations during the backward pass. This allows us to process up to 1.6$\times$ longer visual inputs with just less than 8% overhead in terms of iteration time.
@@
We perform comprehensive evaluation of LV-XAttn on Llama 3-V, mPLUG-Owl3 (Ye et al., 2024) models and OpenFlamingo (Awadalla et al., 2023) models across multiple cluster configurations, including setups with A100 and A30 GPUs. LV-XAttn speeds up the cross-attention operation by up to 45.85$\times$ and overall model iteration time by up to 10.62$\times$ compared to Ring Attention. By minimizing communication volume and further overlapping communication with computation, we demonstrate that LV-XAttn incurs less than 0.42% overhead compared to the theoretical no-communication baseline. LV-XAttn is available at https://github.com/uw-mad-dash/LV-XAttn.
@@ 2 Background
Multimodal Large Language Models As LLMs continue to evolve, researchers are investigating how to incorporate vision and other modalities into these models. One common way is to embed cross-attention layers into the language model. This design has been adopted by a number of models including Flamingo (Alayrac et al., 2022), Otter (Li et al., 2023a), mPLUG-Owl3 (Ye et al., 2024), IDEFICS (Laurençon et al., 2024), Llama 3-V (Grattafiori et al., 2024), NVLM-X (Dai et al., 2024), and NVLM-H (Dai et al., 2024). Broadly, these models follow the architecture illustrated in Figure 1. They include a visual encoder, a projection layer, and an LLM. Cross-attention layers are interleaved between the layers of the LLM. Language inputs are fed directly into the LLM and the resulting intermediate representations are passed to the cross-attention layers as $x$. Visual inputs are processed by the visual encoder and the projection layer to produce visual tokens, which are then passed to the cross-attention layers as $y$, enabling the incorporation of visual information.
@@
Figure 2: Runtime breakdown for a single iteration of Llama 3-V, mPLUG-Owl3-7b, and OpenFlamingo-3b using Ring Attention and LV-XAttn on 16 A100 GPUs. LV-XAttn reduces the time spent on cross-attention computation by 96%, 93%, and 53% for the three models, respectively, compared to Ring Attention. “FWD Vision” refers to the forward pass through the vision encoder and the projection layer; “FWD CA” and “BWD CA” refer to the forward and backward passes through the cross-attention layers in the LLM; and “FWD Non-CA” and “BWD Non-CA” refer to the forward and backward passes through the non-cross-attention layers in the LLM. Llama 3-V was evaluated with a text length of 1K and a frame count of 192 $(S_{Q}=1\text{K},\ S_{KV}=1200\text{K})$; mPLUG-Owl3-7b was evaluated with a text length of 4K and a frame count of 2K $(S_{Q}=4\text{K},\ S_{KV}=1458\text{K})$; and OpenFlamingo-3b was evaluated with a text length and a frame count of 32K $(S_{Q}=32\text{K},S_{KV}=2048\text{K})$.
When applied to scenarios with large visual inputs, cross-attention requires significant memory resources and therefore presents a scaling challenge. The standard implementation of attention involves materializing the matrix product $QK^{T}\in\mathbb{R}^{S_{Q}\times S_{KV}}$, resulting in memory complexity that scales with the product of text sequence and the visual sequence length. The large amount of visual tokens from long videos inputs thus causes a memory bottleneck. For example, in Llama 3-V, a 20-minute video sampled at 1 fps is encoded to a visual input with more than 7 million tokens.
While memory-efficient methods like FlashAttention (Dao et al., 2022; Dao, 2024) reduce the memory footprint of attention operations to enable handling longer context lengths, the amount of memory required still often surpasses the capacity of a single worker. For example, processing a 20-minute long video with a language input of 2048 tokens in Llama 3-V demands over 234 GB of memory for the cross-attention operation, even with FlashAttention.
@@
$O_{i}$ is computed iteratively by transmitting key-value blocks among workers in a ring-like fashion. To facilitate the block-wise computation of $O_{i}$, worker $i$ has to maintain necessary softmax statistics $L_{i}\in\mathbb{R}^{\frac{S_{Q}}{n}}$. During round $r$, worker $i$ computes partial attention using the blocks $Q_{i}$, $K_{(i-r)\mod n}$, and $V_{(i-r)\mod n}$ and updates $O_{i}$ and $L_{i}$. The key-value block is then sent to worker $i+1$ while a new key-value block is received from worker $i-1$.
While Ring Attention can be used to distribute the cross-attention operation, the presence of large key-value blocks makes Ring Attention communication-bound, resulting in highly inefficient cross-attention operations. As illustrated in Figure 2, despite comprising only 3% of the total parameters, cross-attention operations can account for up to 88% of the iteration time when using Ring Attention.
@@ 3.1 Method
         Send $O_{j_{\text{prev}}},L_{j_{\text{prev}}},Q_{j}$ to worker $(i+1)\bmod n$
         Recv $O_{j},L_{j},Q_{j_{\text{next}}}$ from worker $(i-1)\bmod n$
         $\Delta O,\Delta L\leftarrow\text{FlashAttention}(Q_{j},K_{i},V_{i})$
     $O_{j},L_{j}\leftarrow\text{Rescale}(O_{j},L_{j},\Delta O,\Delta L)$
@@
The primary observation that motivates our work is that, in applications involving large visual inputs, the size of the query block is typically much smaller than that of the key-value blocks. For instance, in the widely-used video understanding benchmark Video-MME (Fu et al., 2024), videos have an average duration of 2,386 seconds. Each frame is encoded by the visual encoder and projection layer in an MLLM into multiple visual tokens. For example, Llama 3-V generates 6404 visual tokens per frame. With a sampling rate of 1 fps, each video results in a key-value sequence length of $S_{KV}=15279944$. On the other hand, the average text prompt for long videos consists of $3128$ words, including the question, options, answer, and subtitles, resulting in a query sequence length of $S_{Q}=2386+3128=5514$. As a result, distributed attention mechanisms that involves movement of key-value blocks incur substantial communication overhead.
To address this, we propose LV-XAttn, which keeps the large key-value blocks locally on each worker, while smaller query blocks, attention blocks, and necessary softmax statistics are exchanged among workers in a ring-style fashion. This is illustrated in Figure 3. During each round, each worker $i$ computes attention using its local key-value blocks $K_{i}$ and $V_{i}$ and query blocks $Q_{j}$ received from peers. This computation generates partial attention blocks $\Delta O$ and partial softmax statistics $\Delta L$. The worker then updates the received attention block $O_{j}$ and softmax statistics $L_{j}$ by rescaling them using $\Delta O$ and $\Delta L$. The worker then sends $Q_{j}$, $O_{j}$ and $L_{j}$ to the next worker in the ring topology and receives $Q_{j-1}$, $O_{j-1}$ and $L_{j-1}$ from the previous worker. After $n$ rounds, the computed attention block $O_{i}$ and softmax statistics $L_{i}$ are returned to worker $i$.
@@
  • Receive $O_{j}$ and $L_{j}$ from worker $i-1$, which are needed for rescaling in this round.
@@
  • Send $O_{j+1}$ and $L_{j+1}$ computed in the previous round to worker $i+1$.
@@
After receiving $O_{j}$ and $L_{j}$, and computing $\Delta O$ and $\Delta L$, we can perform rescaling to update $O_{j}$, and $L_{j}$. We describe our distributed attention procedure in Algorithm 1. As demonstrated in Section 4.4, the substantial reduction in communication volume enables complete overlap with computation, effectively eliminating any communication overhead.
Figure 4: The theoretical speedup of LV-XAttn over Ring Attention for cross-attention on a 4-node cluster. Each node is equipped with 4 A100 GPUs, and nodes are interconnected by a 25 GB/s network. The markers represent processing a 2,386-second video and a 3,128-word text prompt – average values for long videos in Video-MME (Fu et al., 2024) – using LLama-3V, mPLUG-Owl3 and OpenFlamingo models at different frame rates. Note that for each frame, a special token <image> have to be added to the text-prompt, resulting in $S_{Q}=2386+3128=5514$.
Table 1: Runtime analysis for Ring Attention and LV-XAttn. $h$ represents the number of heads in multi-head attention. The attention FLOPs are calculated similar to previous work (Dao, 2024).
@@
$\max\left(f(\frac{S_{Q}d}{n},\frac{S_{KV}d}{n}),comm(2\cdot\frac{S_{Q}d}{n}+\frac{S_{Q}}{n})\right)$ (1)
@@ 3.2 Activation Recomputation for MLLM
In standard attention implementation, during forward pass computation, input tensors $Q_{i},K_{i},V_{i}$ and output tensors $O_{i}$ and $L_{i}$ are saved for backward pass. However, storing large key-value blocks $K_{i}$ and $V_{i}$ increases memory usage, thereby limiting the maximum number of visual inputs that can be processed. For instance, in the case of mPLUG-Owl3-7b with a 3600-frame video, storing $K_{i}$ and $V_{i}$ takes $70.08$GB per cross-attention layer.
@@
As demonstrated in the ablation study in Section 4.4, this approach incurs a runtime overhead of less than $8\%$ while enabling the system to handle 1.6×1.6$\times$1.6 × more visual inputs.